40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun May 16 12:51:08 2010, kiwamu, Update, Green Locking, reflected beam at PD 
    Reply  Sun May 16 19:25:45 2010, Koji, Update, Green Locking, reflected beam at PD 
       Reply  Thu May 20 13:29:40 2010, kiwamu, Update, Green Locking, mode profile at PPKTP crystal PPKTPmode.png
          Reply  Thu May 20 14:18:59 2010, kiwamu, Update, Green Locking, mode profile of 40m cavity MMT40mcavity.png
             Reply  Wed May 26 04:14:21 2010, kiwamu, Update, Green Locking, locked OLTF_endPDH.png
                Reply  Wed May 26 14:38:02 2010, Koji, Update, Green Locking, locked 
                   Reply  Thu May 27 02:22:24 2010, kiwamu, Update, Green Locking, more details stimes.pngspectrum.pngETMX_small.png
                      Reply  Thu May 27 09:43:50 2010, rana, Update, Green Locking, more details 
                         Reply  Mon Jul 12 13:16:53 2010, kiwamu, Update, Green Locking, PZT feedback at X end arm_day_night.png
                      Reply  Wed Jun 9 11:13:11 2010, kiwamu, Update, Green Locking, lock with PDH box 
                         Reply  Fri Jun 11 11:54:42 2010, kiwamu, Update, Green Locking, end PDH with thermal feedback LPF.pngthermal_feedback.png
                            Reply  Fri Jun 11 13:32:28 2010, Koji, Update, Green Locking, end PDH with thermal feedback 
                         Reply  Sun Jun 20 18:28:25 2010, kiwamu, Update, Green Locking, Re: lock with PDH box 
                Reply  Wed Jul 21 19:11:18 2010, Dmass, Update, Green Locking, locked 
Message ID: 2936     Entry time: Sun May 16 12:51:08 2010     Reply to this: 2937
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: reflected beam at PD 

Mode matching to the cavity has been done.

Now the reflection from the cavity is successfully going into the PD.

However I could not see any obvious error signal.

I should compute and re-check the expected signal level.

 


(mode matching of the crystal)

On the last Wednesday, Kevin and I measured the mode profile before the PPKTP crystal,  and we found the Gaussian beam at the crystal is focused too tightly (w = 38 um).

In order to achieve the best conversion efficiency the waist size should be 50.0 um. So we moved a lens, which was located before the crystal, to 7 cm more away from the crystal. Eventually we obtained a better focus (w = 50.1 um).

Thanks, Kevin. You did a good job.

 

(mode matching of the cavity)

I put a lens with f=-50 mm after the crystal to diverge the green beam more quickly. Then the beam is going through the Faraday of 532 nm, two final modematching lenses and ETMY at last.

By shifting the positions of these lenses, I obtained the reflection from ITMY with almost the same spot size as that of the incident. This means modemathing is good enough.

I put two more steering mirrors before its injection to the ETM, this allows us to align the beam axis against the cavity.

I aligned the axis by using the steering mirrors and now the green beam are successfully hitting the center of both the ETM and the ITM.

Then the alignment of the ETM and the ITM was adjusted from medm, so that both reflection goes in the same path as that of the incident.

And then I put a PD (Thorlabs PDA36A) to see the reflection rejected by the Faraday.

Connecting a mixer and a local oscillator (Stanford func. generator) with f=200kHz, but I couldn't see any obvious PDH signal....

Since the PD is DC coupled, the signal is almost dominated by DC voltage. Even if I inserted a high pass filter to cut off the DC, the AC signal looks very tiny..

ELOG V3.1.3-